OpenStudy (anonymous):

Mark had 1/2 cup sugar and borrowed 3/4 cup from Cary to add to his. How many cups did he have now? A. 4/6 B. 5/4 C. 1/4 anyone else but mathmathmath

OpenStudy (anonymous):

1/2 + 3/4 = 0.5 + 0.75 = 1.25 5/4 = 1.25 Hence B